### v1

# Investigating Large Language Models in Diagnosing Students’ Cognitive Skills in Math Problem-solving

Abstract
[…]
1 Introduction
Automatic assessment of students’ responses to mathematical problems can support learning at scale. Assessment can serve as feedback to guide students through problem-solving processes, allowing them to refine their reasoning and correct misconceptions as they arise. As teachers’ manual assessment is limited in scale and availability (Stankous, 2016), automated assessment is particularly instrumental in environments where human instructors are unavailable or where student-to-teacher ratios are too high (e.g., MOOCs).
The advancement in AI technologies has expanded the automatic assessment to evaluate more complicated forms of student responses. While early automated assessment methods worked on grading categorical multiple-choice questions, current multimodal and reasoning AI models (e.g., GPT-o1) can process handwritten responses to constructed response questions (Gao et al., 2024; Liu et al., 2024; Chamieh et al., 2024; Baral et al., 2023). As opposed to multiple choices and short answers, these students’ handwritten constructed responses capture students’ thinking processes (Livingston, 2009).
Evaluating students’ constructed responses with AI presents a promising opportunity to diagnose their cognitive skills at scale. Cognitive skills are the abilities to know, apply, and reason mathematical concepts and formulas (Mullis, 2017). Unlike traditional grading that focuses primarily on final answers, cognitive diagnosis enables more granular feedback by identifying where a student’s thinking process may have broken down (Jin et al., 2024; Rahbarnia et al., 2014; Kheong, 1994). For instance, such evaluation can reveal whether a student successfully recalled relevant concepts, represented a problem with an appropriate model, determined a reasonable strategy, and implemented it correctly. These cognitive skills (Mullis, 2017) (complete list in Table 3) often manifest in varied ways through texts and graphics and in subtle ways without overt evidence in students’ responses.
However, such an in-depth diagnosis is challenging because AI models need to understand complicated handwritten responses and need to infer students’ thinking processes beyond visible cues. This diagnosis task involves perceiving graphics in students’ responses, interpreting mathematical notations, and reasoning about students’ problem-solving processes. Existing work has focused on examining AI models’ performance on each of these capabilities in math problem-solving (Didolkar et al., 2024; Fang et al., 2024; Ahn et al., 2024), multimodal perception (Zhang et al., 2024b; a), and reasoning (Wang et al., 2023; Hao et al., 2024) in diverse domains. Whereas previous research has asked how well LLMs can emulate human cognition, our task inverts that perspective, challenging models to infer humans’ logical cognition from partial visual and textual evidence. The diagnosis task is precisely the intersection of these capabilities that makes it particularly complex for current LLMs.
In this work, we investigate how well existing LLMs diagnose students’ cognitive skills in mathematics. Specifically, we address two research questions:
  1. RQ1.
    How accurately do state-of-the-art LLMs diagnose students’ cognitive skills?
  2. RQ2.
    How do multimodal input, reasoning, and model size influence LLMs’ cognitive skill diagnosis?
To answer these questions, we constructed MathCog, a benchmark dataset designed to evaluate cognitive skill diagnosis. In collaboration with 5 education experts and 15 middle school teachers, we curated 12 middle school math problems, each with 50 student responses diagnosed by teachers based on a diagnostic checklist. Each checklist is based on a well-founded diagnosis framework (Mullis, 2017) and consists of 4 to 6 cognitive skill check items (see Table 1).
[…]
  • Creating a benchmark dataset for evaluating cognitive skill diagnosis performance, MathCog, composed of expert-reviewed diagnostic checklists and teacher-diagnosed results for 639 student responses across 12 middle school math problems.
  • Examining the performance of 16 state-of-the-art closed and open LLMs on diagnosing students’ cognitive skills using MathCog.
  • Investigating the impact of multimodal input, reasoning, and model size on the cognitive skill diagnosis task through a comparison study.
2 Related Work
We review the foundational framework underlying our cognitive diagnosis task, followed by prior investigations into LLMs’ performance in math-related contexts, and existing benchmarks, highlighting both their relevance and current gaps.
2.1 Cognitive Diagnosis in Mathematics Education
The TIMSS (Trends in International Mathematics and Science Study) framework (Mullis, 2017) provides a comprehensive and mathematics-specific approach to cognitive diagnosis, comprising content and cognitive domains. The cognitive domain, which evaluates knowledge application, is divided into three key areas: Knowing (recalling definitions, recognizing mathematical entities, computing operations), Applying (determining strategies, representing problems, implementing solutions), and Reasoning (analyzing, integrating, evaluating, justifying). While other frameworks like Bloom’s taxonomy (Bloom et al., 1956) and Polya’s problem-solving framework (Polya, 1957) exist, TIMSS’s mathematics-specific approach, validated through its extensive use across approximately 70 countries since 1995 (Mullis et al., 2023), provides more targeted and reliable guidance for assessing students’ mathematical thinking. Our cognitive skill diagnosis task is founded on this framework that provides deeper insights into students’ mathematical thinking processes, addresses limitations of knowledge-centered assessments, and offers specific guidelines for educational interventions.
2.2 LLM Capabilities in Mathematical Tasks
LLMs have shown significant progress in mathematical problem-solving, with advanced models achieving improved reasoning capabilities through reinforcement learning (Guo et al., 2025; OpenAI, 2024; Zhong et al., 2024). While larger models and Chain-of-Thought prompting further enhance mathematical reasoning capabilities (Li et al., 2024; Yang et al., 2024a; Wei et al., 2022), current models still struggle with multimodal inputs such as visual elements and handwritten content (Zhang et al., 2024b; Baral et al., 2025; Liu et al., 2024). Although domain-specific models like MathGLM-Vision (Yang et al., 2024b) show promise in addressing these challenges, their performance remains limited by data scarcity (Yan et al., 2024). Beyond direct problem-solving, LLMs need metacognitive abilities to evaluate student responses and identify misconceptions (Baral et al., 2024; Kaggle & Eedi, 2020), yet current research focuses mainly on basic error identification rather than complex cognitive diagnosis. We address a critical gap in the field by providing a comprehensive evaluation of LLMs’ ability to track thinking processes and assess cognitive abilities in mathematical contexts.
2.3 Benchmarks for Evaluating Mathematical Tasks
A diverse array of benchmarks exists for evaluating LLMs’ mathematical capabilities, including GSM8K (Cobbe et al., 2021) for elementary school problems, MATH (Hendrycks et al., 2021) for high school competition questions, MathOdyssey (Fang et al., 2024), and DeepMind’s Mathematics dataset (Saxton et al., 2019). In the educational domain, specialized assessments have emerged, such as MathFish (Lucy et al., 2024), which aligns 9,900 problems with 385 K-12 standards to evaluate models’ capacity to recognize specific skills and concepts, along with handwriting recognition (Baral et al., 2025) and misconception detection in MCQs (Kaggle & Eedi, 2020). However, with few exceptions like (Hellman et al., 2023), there is a notable gap in datasets designed for cognitive diagnosis in open-ended responses. While existing research has mainly focused on simple error identification and scoring, the diagnosis of students’ complex cognitive processes in open-ended questions has been relatively less explored. For our novel cognitive skill diagnosis task, we constructed a dataset that supports a comprehensive evaluation of LLMs’ ability to track implicit thinking processes, identify root causes of errors, and assess students’ cognitive abilities.
3 Dataset: MathCog
To evaluate LLMs’ performance on the cognitive skill diagnosis, we first created a benchmark dataset composed of secondary school math problems, student responses, diagnostic checklists, and verdicts (Table 1). The math problem and student response data are from AI-Hub [^1] https://www.aihub.or.kr, which contains 5,932 K-12 math problem types and 147,882 handwritten student response images labeled with OCR transcriptions. Each problem type has 5 to 7 isomorphic problems that have the same problem-solving procedures and differ only in the numbers in the problem description (e.g., “Solve $x^{2}+2x-3=0$” and “Solve $x^{2}+x-2=0$”). Among the data, we chose the problem types and student responses corresponding to grades 7-9, as the problems were complex enough to elicit cognitive processes. Among 2,605 problem types from grades 7 to 9, we excluded the problem types that had fewer than 50 student responses to ensure that we had enough student responses to use in our performance evaluation. Finally, we manually reviewed the remaining 35 problem types and removed the problem types whose corresponding isomorphic problems require different mathematical knowledge (e.g., “Solve $x^{2}+2x-3=0$” (quadratic) vs. “Solve $x^{3}+4x-1=0$” (trinomial)). The filtering process left us with 137 problems and 796 student responses from 15 problem types, each involving the same mathematical knowledge and problem-solving procedures. The specific numbers at each stage of filtering are shown in Figure 1.
Figure 1: The dataset creation process and the number of problem types, problems, and student responses at each filtering stage.
Problem Student Response Diagnostic Checklist Verdict
When $P=\sqrt{3}$ $-7\sqrt{5}+2\sqrt{3}$, $Q=2\sqrt{3}-\sqrt{5}$ $+2\sqrt{5}$, find the value of $P+Q$. [Uncaptioned image]
Recognize: Does the student
realize that if the numbers in
the radicals are the same, he
or she can perform the four
arithmetic operations on the
coefficients with the same
letters?
Vague
Yes
Compute: Were the addition
and subtraction calculations
of coefficients performed
accurately?
Evident
Yes
Determine: Does the student
choose a strategy to first
simply organize each equation
(e.g., put $\sqrt{5}$ terms together,
etc.) and then find the final
sum?
Evident
No
There is a trapezoid with a lower side length of $8$ $cm$ and a height of $4$ $cm$. If the area of this trapezoid is not less than $28$ $cm^{2}$, find how much more $cm$ the length of the upper side of the trapezoid must be. [Uncaptioned image]
Recall: Does the student
remember the formula for
the area of a shape correctly
(e.g., the formula for the area
of a trapezoid)?
Evident
Yes
Compute: Has the student
calculated the linear and
constant terms correctly?
Vague
No
Determine: Did the student
know the need to set up an
equation and then solve it
to find the range of solutions
that meet the conditions?
Evident
Yes
Represent: Has the given
situation been expressed
correctly?
Evident
No
Implement: When simplifying
an expression, does the student
keep the expression correct by
performing the same operation
on both sides?
Vague
Yes
Table 1: Two samples from MathCog. Each data point is composed of a math problem, student response, relevant diagnostic checklist, and verdict for each check item.
3.1 Diagnostic Checklist
For each problem type, we made a diagnostic checklist commonly applicable to isomorphic problems. These checklists are a list of binary question items mapped to one of the 15 cognitive skills defined in the TIMSS 2019 assessment framework (Mullis, 2017). The authors initially drafted the checklist items and refined them by requesting a review from five math curriculum and evaluation experts with PhD degrees in education and practical experience in making math assessment guidelines (e.g., TIMSS, Korean public school curriculum). Experts gave feedback on the clarity, granularity, and validity of the checklist items. For example, the experts suggested changing “square root term” into “irrational term” for generality and combining items like “Were coefficients calculated correctly?” and “Were addition and subtraction performed correctly?” into “Were the addition and subtraction calculations of coefficients performed accurately?” for reasonable granularity: specific enough to describe a distinct skill, yet general enough to accommodate diverse forms of valid evidence. The experts also commented on the skills each problem can or cannot assess. Experts pointed out that our math problems primarily focus on calculating numbers and applying knowledge and, hence, are limited in assessing “reasoning” (e.g., justifying, analyzing, generalizing) (Mullis, 2017) by their design. We scoped our check items to “knowing” and “applying” cognitive domains only. We took two iterations to refine the checklists, and each checklist was reviewed by two experts independently in each iteration.
3.2 Teacher-generated Verdict
We recruited 15 middle school math teachers to evaluate 796 student responses based on predefined diagnostic checklists. These teachers had an average of 6.1±4.3 years of experience (range: 2.5–20 years). Each check item was assessed along two dimensions: “yes/no” and “evident/vague” (see Table 1). A “yes” response indicated that the student fully demonstrated the cognitive action specified, while a “no” indicated otherwise. “Evident” meant there was clear evidence to support the judgment, whereas “vague” signified insufficient evidence. For example, for the check item “Did the student calculate correctly?”, teachers marked “evident yes” if the response showed a complete, error-free calculation process, “evident no” if there were clear mistakes, and “vague no” if calculations were missing.
To account for the subjectivity in the diagnosis, we grouped three teachers to evaluate approximately 160 student responses together. The diagnosis occurred in two phases: first, each group collaboratively assessed 36 responses, resolving conflicts through majority voting; second, each teacher independently assessed 18 responses. The second phase included 18 overlapping responses, allowing us to measure inter-rater agreement (see Table  4 in Appendix). Based on the threshold established in prior literature (Graham et al., 2012), we set a minimum agreement of 70% and excluded data from problem types 6, 13, and 14 accordingly. The distribution of the verdicts is presented in Table 5.
4 Experimental Setting
Using MathCog as a benchmark dataset, we explored two research questions. For RQ1, we evaluated the state-of-the-art LLMs, testing GPT (Achiam et al., 2023), Gemini (Team et al., 2023), Llama (Touvron et al., 2023), Claude (Anthropic, 2024), and DeepSeek (Guo et al., 2025) series. For RQ2, we configured three experiments in which we compared the performance of LLMs in different inputs, reasoning capabilities, and model sizes.
Task Definition. Our task for LLMs is essentially a single-label classification task. Formally, we represent diagnosis result for a student’s response as a collection of tuples $(c,r)$, where $c\in C$ (i.e., check items) and $r\in\{Evident\ Yes,Vague\ Yes,Evident\ No,Vague\ No\}$. The combined judgment $(c,r)$ provides both an evaluative decision and a confidence signal for a skill. Given a context $(P,S,R,C)$, where P is the problem, S is the answer, and R is the student response, the task is to produce assessment tuples $(c,r)$ for all $c\in C$.
Prompting. We instructed LLMs to evaluate each check item based on a given math problem, its solution, a student’s response, and a diagnostic checklist. The LLMs processed inputs in the form of OCR transcriptions, where mathematical formulas and visual cues (e.g., strikethroughs) were represented using LaTeX. In the case of testing the effect of multimodality, we supplied images of student responses. Since the original inputs were in Korean, we machine-translated them into English to prevent possible performance degradation due to language (Achiam et al., 2023). To enhance response accuracy, we employed Chain-of-Thought prompting (Wei et al., 2022; Liu et al., 2024), guiding the LLM to systematically address each check item by first restating its content, identifying relevant evidence, providing an explanation, and delivering a final verdict. The verdict followed one of the four categories used by teachers in MathCog. We opted not to include few-shot examples, as MathCog does not provide explicit evidence and explanations for reference. To minimize randomness, all LLMs were run with a temperature setting of zero. A detailed description of the system and user prompts can be found in the Appendix.
Models. We conducted our experiment using 16 distinct closed and open LLMs from various vendors. To assess the impact of multimodal input, we tested GPT-4o (2024-08-06), Claude-3.5-Sonnet (20240620), and Gemini-1.5-Flash under two conditions: one using text-only prompts and another supplemented with an image. To examine the effect of reasoning capabilities, we compared LLMs with reasoning capabilities—GPT-o1-Preview (2024-09-12), Gemini-2.0-Flash-Thinking (exp-01-21), and DeepSeek-R1—with conventional LLMs of the same model series—GPT-4o, GPT-4o-Mini (2024-07-18), Gemini-1.5-Pro (002), Gemini-1.5-Flash, and DeepSeek-V3 (0324). Finally, we investigated model size differences: GPT-4o, Gemini-1.5-Pro, Llama-3.1-450B (2024-07-23), and DeepSeek-V3 were categorized as large models; Llama-3.3-70B (2024-07-23) and Qwen-2.5-72B (2024-09) as medium models; and GPT-4o-Mini, Gemini-1.5-Flash-8B (001), Llama-3.1-8B (2024-07-23), Qwen-2.5-7B (2024-09), and Mistral-7B (2024-08-22) as small models.
Metrics. We evaluated the outputs of the large language models (LLMs) against ground-truth diagnostic results provided by human experts. Performance was assessed using both macro F1 score and accuracy computed over verdicts (i.e., $(c,r)$) to quantify the effectiveness of skill diagnosis. For a more nuanced analysis, we also examined the models’ tendencies toward overconfidence and underconfidence. Specifically, overconfidence is defined as the conditional proportion of “evident” verdicts produced by the LLMs that were incorrect. Conversely, underconfidence is the conditional proportion of verdicts labeled as “vague” by the LLMs that were, in fact, both correct and should have been classified as “evident.”
5 Results
In this section, we organize and interpret the findings to answer our research questions. The complete results from our experiment are presented in Table 6 in the Appendix.
5.1 RQ1. How accurately do state-of-the-art LLMs diagnose students’ cognitive skills?
Overall, the results underscore the difficulty of the task (Table 2); no model surpassed an F1 score of 0.5, indicating that even state-of-the-art models struggle to diagnose students’ cognitive skills accurately. Our manual error analysis on a few samples revealed that LLMs largely failed when 1) student response flows in a non-directional, scattered layout, 2) a student uses underscores, strikethroughs, and calculation notes, and 3) a response lacks direct evidence for a skill, despite contextual information being sufficient for inference.
[…]
Models Precision Recall F1 Accuracy Overconfidence Underconfidence
DeepSeek-V3 .406 .461 .417 .714 .818 .107
GPT-4o .398 .476 .412 .672 .788 .110
Llama-3.1-405B .415 .438 .384 .624 .778 .186
Claude-3.5-Sonnet .329 .387 .332 .654 .812 .173
Gemini-1.5-Pro .430 .496 .441 .707 .840 .132
Table 2: The performance of the five state-of-the-art LLMs on the cognitive skill diagnosis task.
Analysis of skill-specific performance (Figure 2) reveals further insights into the limitations of current LLMs. No skill category achieved a maximum F1 score above 0.5. Interestingly, the skill Compute, which most closely resembles conventional grading tasks (i.e., evaluating the correctness of mathematical procedures and calculations), still yielded only a moderate F1 score across models. This moderate F1 score confirms LLMs’ struggle in verifying correct computation in constructed student responses.
Contrary to expectations, there was no substantial performance gap between the “Knowing” and “Applying” cognitive skills. Although we anticipated that “Knowing” skills (e.g., Recall, Recognize) would be easier than “Applying” skills for LLMs due to their surface-level nature, models performed comparably. An exception is found in the skill Classify/Order, which shows a relatively high average accuracy but a lower F1 score. Closer inspection reveals that “Evident Yes” is dominant in the ground truth dataset (Table 5), leading to the discrepancy between F1 score and accuracy.
Figure 2: The performance of five state-of-the-art LLMs across cognitive skill types. The left plot shows the minimum (bottom $\times$), maximum (top $\times$), and average (middle •) F1 scores for each skill category, while the right plot presents the corresponding accuracy metrics.
5.2 RQ2. How do multimodal input, reasoning, and model size influence LLMs’ cognitive skill diagnosis?
We observed subtle impacts of multimodal input and reasoning on the performance, while the impact of model size was significant.
Multimodal Input. Models supporting image input (e.g., GPT-4o-img, Claude-3.5-Sonnet-img, Gemini-1.5-Flash-img) consistently outperformed their text-only counterparts in accuracy (up to 10.6%) and in reducing underconfidence (up to 4.9%) (Figure 3 left), although the statistical difference between the pairs was not significant. Image input was particularly beneficial in cases where OCR transcription introduced errors. For example, a student underlined an arithmetic expression and wrote the corresponding calculation beneath it (see Figure 5 a). The OCR system misinterpreted this layout as a fraction, leading Claude-3.5-Sonnet (text-only) to incorrectly conclude that the student had set up an invalid equation. In contrast, the same model with image input correctly inferred the student’s intention and produced an accurate diagnosis.
Reasoning. The performance of reasoning-oriented models showed mixed results (Figure 3 right). DeepSeek-R1 achieved the highest F1 and accuracy scores among all models. However, other reasoning models, such as o1-Preview and Gemini-2.0-Flash-Thinking, did not show clear advantages over standard models in either accuracy or F1 score. These results suggest that reasoning capability alone does not guarantee better performance. For instance (Figure 5 b), DeepSeek-R1 classified a response as “Vague No” because the student omitted the factorization steps. However, the original expression was $\sqrt{48a}$, and the model failed to detect a clear error in the variable term, indicating that the student’s reasoning was evidently flawed. Throughout its reasoning process, the model focused on the constant term only and overlooked other possible places of error. While current reasoning models excel at pushing forward linear, step-by-step logic, effective cognitive diagnosis may also require divergent, multi-perspective reasoning for thorough analysis.
[…]
Figure 3: Impact of multimodal input (left) and reasoning capability (right) on LLMs’ performance in cognitive skill diagnosis. On the left, each adjacent pair of bars shows performance with and without the image input from the same model family. On the right, models with reasoning capabilities are grouped at the front.
Figure 4: Impact of model size on LLM performance in cognitive skill diagnosis. Models are grouped into large, medium, and small categories.
6 Discussion
Our findings imply that current LLMs are not yet suitable for high-stakes cognitive skill diagnosis. The models exhibit low recall and high overconfidence (Table 2), raising concerns that LLM-based systems may frequently overlook students’ misconceptions. This shortcoming is particularly problematic given that the primary goal of automated assessment is to identify and address such misconceptions, and it can compromise teachers’ and students’ trust in automated systems (Shin, 2021). To responsibly integrate LLMs into real-world assessment settings, practitioners should begin by evaluating their performance on small, ground-truth samples. This process not only exposes limitations but also allows for fine-tuning model judgments and confidence levels for downstream math topics and student populations. In parallel, model developers may work to mitigate overconfidence during training by encouraging more conservative, evidence-based decision-making.
Our manual analysis of failed cases reveals that inferring students’ thought processes from handwritten responses presents distinct challenges. While mathematical problem-solving is inherently procedural, students often express their reasoning in non-linear, spatially scattered layouts that lack clear directionality. This ill-structured presentation differs significantly from the format of most existing reasoning and problem-solving benchmarks. Moreover, ambiguous notations (e.g., fractions misinterpreted as underscores) and unconventional layouts frequently result in OCR transcription errors, which hinder accurate diagnosis unless image inputs are provided. To enable more reliable inference, future work may require richer representations of math responses that capture high-level semantic intent beyond surface-level positional and textual cues.
7 Limitations and Future Work
While MathCogrepresents a novel benchmark for evaluating LLMs on cognitive skill diagnosis, it has several limitations. First, due to data-sharing restrictions from the source provider (AIHub), the original math problems and student responses cannot be publicly released outside Korea. To support future research despite this constraint, we release the complete set of diagnostic checklists and expert-generated human annotations, which can serve as references for researchers to develop similar datasets in their educational contexts. In parallel, we are actively collecting new student response data under informed consent, with the goal of constructing a fully public benchmark in the future.
This ongoing data collection effort will also address limitations in the diversity of problem types represented in MathCog. Currently, the dataset focuses on nine problem types, primarily centered around arithmetic equation solving. Other areas of mathematics—such as geometry—often require interpreting visual representations and may further benefit from the integration of multimodal input. Additionally, reasoning-oriented problems, which explicitly ask students to generalize mathematical knowledge or justify their arguments, are underrepresented. We plan to expand the dataset to include a broader range of content areas and grade levels, enabling a more holistic evaluation of LLM performance across the mathematics curriculum.
Our experimental results are also bound to zero-shot settings, without exploring the effects of few-shot or test-time compute (Snell et al., 2024) prompting. Future work can investigate whether more sophisticated prompting enhances diagnostic accuracy. For example, prompting an LLM to explicitly infer a student’s thinking process before answering each check item may lead to more accurate judgments. We see our work as a first step toward automated cognitive skill assessment, and we invite the community to build upon these directions.
Acknowledgments
This work was supported by Algorithm LABS. This research (paper) used datasets from “The Open AI Dataset Project (AI-Hub, S. Korea).” All data information can be accessed through “AI-Hub (www.aihub.or.kr).”
Appendix A Appendix
Domains Cognitive Skills Description
Knowing Recall
Recall definitions, terminology, number properties, units of measurement,
geometric properties, and notation (e.g., $a\times b=ab,a+a+a=3a$).
Recognize
Recognize numbers, expressions, quantities, and shapes. Recognize entities
that are mathematically equivalent (e.g., equivalent familiar fractions,
decimals, and percents; different orientations of simple geometric figures).
Classify/Order Classify numbers, expressions, quantities, and shapes by common properties.
Compute
Carry out algorithmic procedures for +,,×,÷+,-$\times$,$\div$+ , - , × , ÷ or a combination of these
with whole numbers, fractions, decimals, and integers. Carry out
straightforward algebraic procedures.
Retrieve Retrieve information from graphs, tables, texts, or other sources.
Measure Use measuring instruments; and choose appropriate units of measurement.
Applying Determine
Determine efficient/appropriate operations, strategies, and tools for solving
problems for which there are commonly used methods of solution.
Represent/Model
Display data in tables or graphs; create equations, inequalities, geometric
figures, or diagrams that model problem situations; and generate equivalent
representations for a given mathematical entity or relationship.
Implement
Implement strategies and operations to solve problems involving familiar
mathematical concepts and procedures.
Reasoning Analyze
Determine, describe, or use relationships among numbers, expressions,
quantities, and shapes.
Integrate/Synthesize
Link different elements of knowledge, related representations, and
procedures to solve problems.
Evaluate Evaluate alternative problem solving strategies and solutions.
Draw conclusions Make valid inferences on the basis of information and evidence.
Generalize
Make statements that represent relationships in more general and more
widely applicable terms.
Justify Provide mathematical arguments to support a strategy or solution.
Table 3: Fifteen cognitive skills and their descriptions defined in the TIMSS 2019 framework.
Problem Types 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
% Agreement 95 70 96 95 89 68 74 86 80 80 100 88 40 57 78
Table 4: The inter-rater percentage absolute agreement of each problem type. The percentage indicates the ratio of unanimous verdicts in each teacher group.
A.1 Prompts
The blue text represents the programmatically filled arguments, and the orange text represents LLM-generated output.
A.2 System Prompt
# **Task Description**
You are a middle school math teacher tasked with evaluating students mathematical thinking skills based on their responses to math problems. Your goal is to analyze a given students response and determine whether they exhibit specific cognitive skills in solving the problem. Your evaluation must be **strict** and **evidence-based**, meaning that every verdict must be backed by direct evidence from the response. If no clear evidence exists, do not assume correctness.
## **Evaluation Categories**
For each thinking skill in the checklist, you must classify the students performance into one of the following categories:
- **Evident Yes**: The students response provides clear and explicit evidence that the check item is met. A direct quote from the response can confirm this.
- **Vague Yes**: The response suggests that the check item might be satisfied, but no specific part of the response directly proves it.
- **Evident No**: The response explicitly contradicts or fails to meet the check item, with clear evidence demonstrating the error or omission.
- **Vague No**: The response does not appear to satisfy the check item, but there is no direct evidence confirming whether the student considered it or not.
## **Input Format**
You will receive the following data:
- **Problem**: A math problem given to a student.
- **Answer**: The correct step-by-step solution.
- **Response**: The students response to the problem.
- **Check Items**: A set of specific skills to evaluate.
## **Output Format**
Return a valid JSON object structured as follows:
‘‘‘json
"skills": [
"checkItem": "<Check Items [Label] and the Following Question>",
"evidence": "<Directly Quoted Part of Response>",
"explanation": "<Explanation About Why the Evidence Supports the Verdict>",
"verdict": "Evident Yes" | "Vague Yes" | "Evident No" | "Vague No"
A.3 User Prompt
# **Task**
student responses to math problems, extract direct evidence, and strictly classify thinking skills according to the given categories.
Return a valid JSON object structured as follows:
‘‘‘json
"skills": [
"checkItem": "<Check Items [Label] and the Following Question>",
"evidence": "<Directly Quoted Part of Response>",
"explanation": "<Explanation About Why the Evidence Supports the Verdict>",
"verdict": "Evident Yes" | "Vague Yes" | "Evident No" | "Vague No"
**Problem:**
There is a two-digit natural number whose tens digit is $1$. If the number that changes the tens and ones digits of this natural number is $9$ less than $5$ times the first number, find the first number.
**Answer:**
If the number in the ones place is $x$, this natural number is $10+x$. The number where the tens digit and the ones digit are swapped is $110x+$ 1 because the tens digit is $x$ and the ones digit is $1$. The changed number is $9$ smaller than $5$ times the first number, so $910x+1=5(10+x)-$ 9 $4110x+1=5x+$ 41 $405x=$ 40 $x=8$ Therefore, the first number is $18$.
**Response:**
$1$ $x$ $x$ $1$ $910x+1=5(10+x)-$ 9 $910x+1=50+5x-$ 9 $41$ $405x=$ 40 $x=8$
**Check Items:**
- [recognize_0] Is the student aware that the relationship between two numbers switched can be expressed as addition/subtraction/multiplication?
- [recognize_1] Has the student figured out all the elements necessary to solve the problem (original numbers, swapped numbers, difference or multiple relationships, etc.)?
- [compute_0] Has the student performed algebraic calculations correctly, such as expanding expressions and organizing similar terms?
- [determine_0] Has the student chosen the appropriate approach to solving the problem (e.g., establishing a linear equation)?
- [represent_0] Are two-digit numbers expressed correctly in the form 10x + y (or equivalent)?
- [implement_0] Does the student proceed consistently with the process of expanding the equation and moving terms to find x (or y) according to the planned solution strategy?
**Output:**
‘‘‘json
"skills": [
"checkItem": "[recognize_0] Is the student aware that the relationship between two numbers switched can be expressed as addition/subtraction/multiplication?",
"evidence": "10 x+1=5(10+x)-9",
"explanation": "The student sets up the equation 10x+1=5(10+x)-9, which shows they understand the relationship between the original and swapped numbers involves subtraction and multiplication.",
"verdict": "Evident Yes"
"checkItem": "[recognize_1] Has the student figured out all the elements necessary to solve the problem (original numbers, swapped numbers, difference or multiple relationships, etc.)?",
"evidence": "10 x+1=5(10+x)-9",
"explanation": "The student uses the correct equation to represent the relationship between the original and swapped numbers, indicating they have identified the necessary elements.",
"verdict": "Evident Yes"
Recall Recognize Classify/Order Compute Retrieve Determine Represent Implement
Evident Yes 507 506 92 584 9 441 154 319
Evident No 78 69 8 160 9 54 55 146
Vague Yes 58 54 2 48 0 25 6 29
Vague No 98 62 0 58 34 16 46 41
Student responses 582 583 51 796 52 482 261 535
Table 5: Distribution of verdicts and number of diagnosed student responses in each cognitive skill. Note that the number of student responses can be larger than the sum of the four labels because some diagnostic checklists have two items for the same cognitive skill.
Models Precision Recall F1 Accuracy Overconfidence Underconfidence
DeepSeek-R1 .447 .467 .442 .773 .901 .028
DeepSeek-V3 .406 .461 .417 .714 .710 .107
GPT-4o-img .427 .494 .448 .743 .738 .070
GPT-4o .398 .476 .412 .672 .703 .110
GPT-4o-Mini .345 .397 .352 .622 .791 .102
o1-Preview .338 .410 .354 .711 .757 .117
Llama-3.3-70B .416 .468 .400 .640 .534 .189
Llama-3.1-8B .283 .265 .257 .642 .695 .138
Llama-3.1-405B .415 .438 .384 .624 .451 .186
Claude-3.5-Sonnet-img .326 .365 .330 .690 .680 .140
Claude-3.5-Sonnet .329 .387 .332 .654 .676 .173
Gemini-1.5-Flash-img .431 .484 .429 .702 .593 .129
Gemini-1.5-Flash-8b .287 .341 .279 .558 .579 .196
Gemini-1.5-Flash .432 .490 .432 .679 .582 .155
Gemini-1.5-Pro .430 .496 .441 .707 .721 .132
Gemini-2.0-Flash-Thinking .455 .492 .421 .690 .589 .186
Mistral-7B .384 .369 .347 .714 .805 .112
Qwen-2.5-7B .278 .270 .271 .705 .842 .075
Qwen-2.5-72B .415 .459 .416 .710 .631 .115
Table 6: Performance comparison of all 16 LLMs tested on skill diagnosis tasks.
Figure 5: Illustrative examples of diagnosis check items and student responses that LLMs failed to diagnose correctly. Evidence for human judgment is marked with a red box.

### v2

@@
-Investigating Large Language Models in Diagnosing Students’ Cognitive Skills in Math Problem-solving
+Benchmarking Large Language Models for Diagnosing Students’ Cognitive Skills from Handwritten Math Work
 Abstract
-[…]
+Students’ handwritten math work provides a rich resource for diagnosing cognitive skills, as it captures intermediate reasoning beyond final answers.
+We investigate how current large language models (LLMs) perform in diagnosing cognitive skills from such work.
+However, student responses vary widely, often omitting steps or providing only vague, contextually implicit evidence.
+Despite recent advances in LLMs’ multimodal and reasoning capabilities, their performance under such conditions remains underexplored.
+To address this gap, we constructed MathCog, a benchmark dataset containing 3,036 diagnostic verdicts across 639 student responses to 110 math problems, annotated by teachers using TIMSS-grounded cognitive skill checklists with evidential strength labels (Evident/Vague).
+Evaluating 18 LLMs, we find that (1) all models underperform (F1 < 0.5) regardless of capability, and (2) performance degrades sharply under vague evidence.
+Error analysis reveals systematic patterns: models frequently misattribute vague evidence as evident, overthink minimal cues, and hallucinate nonexistent evidence.
+We discuss implications for evidence-aware, teacher-in-the-loop designs for LLM-based cognitive diagnosis in educational settings.
 1 Introduction
-Automatic assessment of students’ responses to mathematical problems can support learning at scale.
-Assessment can serve as feedback to guide students through problem-solving processes, allowing them to refine their reasoning and correct misconceptions as they arise.
-As teachers’ manual assessment is limited in scale and availability (Stankous, 2016), automated assessment is particularly instrumental in environments where human instructors are unavailable or where student-to-teacher ratios are too high (e.g., MOOCs).
-The advancement in AI technologies has expanded the automatic assessment to evaluate more complicated forms of student responses.
-While early automated assessment methods worked on grading categorical multiple-choice questions, current multimodal and reasoning AI models (e.g., GPT-o1) can process handwritten responses to constructed response questions (Gao et al., 2024; Liu et al., 2024; Chamieh et al., 2024; Baral et al., 2023).
-As opposed to multiple choices and short answers, these students’ handwritten constructed responses capture students’ thinking processes (Livingston, 2009).
-Evaluating students’ constructed responses with AI presents a promising opportunity to diagnose their cognitive skills at scale.
-Cognitive skills are the abilities to know, apply, and reason mathematical concepts and formulas (Mullis, 2017).
-Unlike traditional grading that focuses primarily on final answers, cognitive diagnosis enables more granular feedback by identifying where a student’s thinking process may have broken down (Jin et al., 2024; Rahbarnia et al., 2014; Kheong, 1994).
-For instance, such evaluation can reveal whether a student successfully recalled relevant concepts, represented a problem with an appropriate model, determined a reasonable strategy, and implemented it correctly.
-These cognitive skills (Mullis, 2017) (complete list in Table 3) often manifest in varied ways through texts and graphics and in subtle ways without overt evidence in students’ responses.
-However, such an in-depth diagnosis is challenging because AI models need to understand complicated handwritten responses and need to infer students’ thinking processes beyond visible cues.
-This diagnosis task involves perceiving graphics in students’ responses, interpreting mathematical notations, and reasoning about students’ problem-solving processes.
-Existing work has focused on examining AI models’ performance on each of these capabilities in math problem-solving (Didolkar et al., 2024; Fang et al., 2024; Ahn et al., 2024), multimodal perception (Zhang et al., 2024b; a), and reasoning (Wang et al., 2023; Hao et al., 2024) in diverse domains.
-Whereas previous research has asked how well LLMs can emulate human cognition, our task inverts that perspective, challenging models to infer humans’ logical cognition from partial visual and textual evidence.
-The diagnosis task is precisely the intersection of these capabilities that makes it particularly complex for current LLMs.
-In this work, we investigate how well existing LLMs diagnose students’ cognitive skills in mathematics. Specifically, we address two research questions:
-  1. RQ1.
-    How accurately do state-of-the-art LLMs diagnose students’ cognitive skills?
-  2. RQ2.
-    How do multimodal input, reasoning, and model size influence LLMs’ cognitive skill diagnosis?
-To answer these questions, we constructed MathCog, a benchmark dataset designed to evaluate cognitive skill diagnosis.
-In collaboration with 5 education experts and 15 middle school teachers, we curated 12 middle school math problems, each with 50 student responses diagnosed by teachers based on a diagnostic checklist.
-Each checklist is based on a well-founded diagnosis framework (Mullis, 2017) and consists of 4 to 6 cognitive skill check items (see Table 1).
-[…]
-  • Creating a benchmark dataset for evaluating cognitive skill diagnosis performance, MathCog, composed of expert-reviewed diagnostic checklists and teacher-diagnosed results for 639 student responses across 12 middle school math problems.
-  • Examining the performance of 16 state-of-the-art closed and open LLMs on diagnosing students’ cognitive skills using MathCog.
-  • Investigating the impact of multimodal input, reasoning, and model size on the cognitive skill diagnosis task through a comparison study.
+Diagnosing students’ cognitive skills from their problem-solving work is a long-standing goal in mathematics education, as it can reveal where and how students’ reasoning breaks down beyond final correctness [17, 12].
+As defined in the TIMSS assessment framework [22], cognitive skills refer to students’ abilities to know, apply, and reason with mathematical concepts and procedures.
+Handwritten math work is particularly valuable for diagnosing these skills because it can reflect intermediate reasoning steps and partial understanding [14, 12].
+However, such diagnostic evidence is often incomplete, implicit, or unevenly expressed across students [11, 24], making reliable cognitive skill diagnosis inherently challenging.
+Despite recent advances in large language models (LLMs), including multimodal perception [31, 30] and reasoning capabilities [27, 9], their ability to diagnose students’ cognitive skills from handwritten work remains largely unexamined.
+While prior work has examined LLMs’ own mathematical problem-solving abilities [5, 6, 2], our work extends beyond problem solving to examine their ability to diagnose human problem-solving processes.
+In this work, we systematically investigate how well existing LLMs diagnose students’ cognitive skills in mathematics.
+Specifically, we focus on the degree of evidential strength in students’ handwritten responses.
+We address the following research questions:
+  • RQ1.
+    How do different LLMs (varying in image input, reasoning, model size, and few-shot prompting) perform in diagnosing students’ cognitive skills?
+  • RQ2.
+    How does the evidential strength of student responses affect LLM’s cognitive skill diagnosis performance?
+To answer these questions, we constructed MathCog, an expert-crafted benchmark dataset designed to evaluate cognitive skill diagnosis.
+In collaboration with 5 education experts and 15 middle school teachers, we curated 12 middle school math topics and 110 problems, each with 50+ student responses diagnosed by teachers based on a problem-specific diagnostic checklist grounded in the TIMSS cognitive framework [22].
+Teachers provided binary judgements (Yes/No) for each skill and annotated the presence of supporting evidence as Evident or Vague, yielding 3,036 diagnostic items total.
+Critically, we curated our dataset to include only problems where human experts achieve >70% agreement, establishing a stable ground truth rather than subjective noise.
+Using MathCog, we evaluated 18 closed- and open-source LLMs spanning multiple model families, sizes, and capabilities.
+Results indicate that current LLMs struggle to reliably diagnose students’ cognitive skills (all F1 scores < 0.5), with performance degrading sharply under vague-evidence conditions and frequently misattributing weak evidence as strong.
+Qualitative error analysis further reveals the model’s failure patterns across model types.
+In particular, models often misidentify or hallucinate supporting evidence, or over-infer students’ cognitive skills from incomplete work, showing that errors cascade throughout the diagnostic pipeline from recognizing student responses to interpreting them based on the rubric.
+Building on these findings, we discuss implications for designing LLM-based diagnostic tools that better elicit students’ diagnostic evidence of cognitive processes and support teacher-in-the-loop interpretation.
+Our contributions are threefold.
+(1) We introduce MathCog, an expert-crafted benchmark dataset for cognitive skill diagnosis.
+(2) Using MathCog, we evaluate 18 large language models and analyze the effects of multimodality, reasoning, model size, and evidential strength on diagnosis performance.
+(3) We analyze diagnostic error patterns under vague-evidence conditions and discuss implications for the design and use of LLM-based cognitive diagnostic systems.
 2 Related Work
-We review the foundational framework underlying our cognitive diagnosis task, followed by prior investigations into LLMs’ performance in math-related contexts, and existing benchmarks, highlighting both their relevance and current gaps.
+We review the foundational framework underlying our cognitive diagnosis task, prior investigations into LLMs’ performance in math-related contexts, and existing benchmarks, highlighting their current gaps.
 2.1 Cognitive Diagnosis in Mathematics Education
-The TIMSS (Trends in International Mathematics and Science Study) framework (Mullis, 2017) provides a comprehensive and mathematics-specific approach to cognitive diagnosis, comprising content and cognitive domains.
-The cognitive domain, which evaluates knowledge application, is divided into three key areas: Knowing (recalling definitions, recognizing mathematical entities, computing operations), Applying (determining strategies, representing problems, implementing solutions), and Reasoning (analyzing, integrating, evaluating, justifying).
-While other frameworks like Bloom’s taxonomy (Bloom et al., 1956) and Polya’s problem-solving framework (Polya, 1957) exist, TIMSS’s mathematics-specific approach, validated through its extensive use across approximately 70 countries since 1995 (Mullis et al., 2023), provides more targeted and reliable guidance for assessing students’ mathematical thinking.
-Our cognitive skill diagnosis task is founded on this framework that provides deeper insights into students’ mathematical thinking processes, addresses limitations of knowledge-centered assessments, and offers specific guidelines for educational interventions.
+The TIMSS [22] is a comprehensive and math-specific framework for cognitive diagnosis, comprising content and cognitive domains.
+The cognitive domain, which evaluates knowledge application, is divided into three key areas: Knowing (recalling definitions, recognizing mathematical entities, classifying or ordering quantities, computing), Applying (determining strategies, representing problems, implementing solution procedures), and Reasoning (analyzing, justifying).
+This research aims to explore whether LLMs can substitute for the mapping to enable scalable and explainable cognitive diagnosis.
+To our knowledge, this is the first systematic investigation of LLMs in the context of the TIMSS framework.
 2.2 LLM Capabilities in Mathematical Tasks
-LLMs have shown significant progress in mathematical problem-solving, with advanced models achieving improved reasoning capabilities through reinforcement learning (Guo et al., 2025; OpenAI, 2024; Zhong et al., 2024).
-While larger models and Chain-of-Thought prompting further enhance mathematical reasoning capabilities (Li et al., 2024; Yang et al., 2024a; Wei et al., 2022), current models still struggle with multimodal inputs such as visual elements and handwritten content (Zhang et al., 2024b; Baral et al., 2025; Liu et al., 2024).
-Although domain-specific models like MathGLM-Vision (Yang et al., 2024b) show promise in addressing these challenges, their performance remains limited by data scarcity (Yan et al., 2024).
-Beyond direct problem-solving, LLMs need metacognitive abilities to evaluate student responses and identify misconceptions (Baral et al., 2024; Kaggle & Eedi, 2020), yet current research focuses mainly on basic error identification rather than complex cognitive diagnosis.
-We address a critical gap in the field by providing a comprehensive evaluation of LLMs’ ability to track thinking processes and assess cognitive abilities in mathematical contexts.
-2.3 Benchmarks for Evaluating Mathematical Tasks
-A diverse array of benchmarks exists for evaluating LLMs’ mathematical capabilities, including GSM8K (Cobbe et al., 2021) for elementary school problems, MATH (Hendrycks et al., 2021) for high school competition questions, MathOdyssey (Fang et al., 2024), and DeepMind’s Mathematics dataset (Saxton et al., 2019).
-In the educational domain, specialized assessments have emerged, such as MathFish (Lucy et al., 2024), which aligns 9,900 problems with 385 K-12 standards to evaluate models’ capacity to recognize specific skills and concepts, along with handwriting recognition (Baral et al., 2025) and misconception detection in MCQs (Kaggle & Eedi, 2020).
-However, with few exceptions like (Hellman et al., 2023), there is a notable gap in datasets designed for cognitive diagnosis in open-ended responses.
-While existing research has mainly focused on simple error identification and scoring, the diagnosis of students’ complex cognitive processes in open-ended questions has been relatively less explored.
-For our novel cognitive skill diagnosis task, we constructed a dataset that supports a comprehensive evaluation of LLMs’ ability to track implicit thinking processes, identify root causes of errors, and assess students’ cognitive abilities.
+Recent advances have enabled LLMs to achieve remarkable performance in mathematical problem-solving [8, 23], yet they show relative weaknesses in diagnosing student abilities [20, 29].
+These diagnostic challenges are further complicated by current models’ struggles with multimodal inputs such as visual elements and handwritten content [31, 4, 16].
+While some recent work has demonstrated LLMs’ potential for cross-domain cognitive diagnosis [19], existing research has limited systematic exploration of LLMs’ diagnostic capabilities for cognitive skills in math specifically.
+We address this critical gap in the field by providing the first comprehensive evaluation of LLMs’ ability to diagnose cognitive skills in mathematical contexts using an established educational framework.
+2.3 Benchmarks for Evaluating Mathematical Assessment Tasks
+In educational settings, several benchmarks have been proposed to evaluate LLMs’ performance in mathematical assessment tasks, including grading, skill recognition, and error analysis.
+For example, MathFish [18] aligns 9,900 problems with 385 K–12 standards to assess models’ ability to identify targeted mathematical skills and concepts.
+Automated scoring of open-ended constructed responses in math word problems has also been studied [10].
+Other benchmarks focus on misconception detection in multiple-choice questions [13] or handwriting recognition in student work [4].
+While prior research has focused on error detection and scoring, diagnosing complex cognitive skills from students’ open-ended handwritten responses remains underexplored and under-resourced.
+To address this gap, we introduce a novel dataset for cognitive skill diagnosis that enables systematic evaluation of LLMs’ ability to interpret implicit reasoning and diagnose students’ cognitive skills beyond final answers.
 3 Dataset: MathCog
-To evaluate LLMs’ performance on the cognitive skill diagnosis, we first created a benchmark dataset composed of secondary school math problems, student responses, diagnostic checklists, and verdicts (Table 1).
-The math problem and student response data are from AI-Hub [^1] https://www.aihub.or.kr, which contains 5,932 K-12 math problem types and 147,882 handwritten student response images labeled with OCR transcriptions.
-Each problem type has 5 to 7 isomorphic problems that have the same problem-solving procedures and differ only in the numbers in the problem description (e.g., “Solve $x^{2}+2x-3=0$” and “Solve $x^{2}+x-2=0$”).
-Among the data, we chose the problem types and student responses corresponding to grades 7-9, as the problems were complex enough to elicit cognitive processes.
-Among 2,605 problem types from grades 7 to 9, we excluded the problem types that had fewer than 50 student responses to ensure that we had enough student responses to use in our performance evaluation.
-Finally, we manually reviewed the remaining 35 problem types and removed the problem types whose corresponding isomorphic problems require different mathematical knowledge (e.g., “Solve $x^{2}+2x-3=0$” (quadratic) vs.
-“Solve $x^{3}+4x-1=0$” (trinomial)).
-The filtering process left us with 137 problems and 796 student responses from 15 problem types, each involving the same mathematical knowledge and problem-solving procedures.
-The specific numbers at each stage of filtering are shown in Figure 1.
-Figure 1: The dataset creation process and the number of problem types, problems, and student responses at each filtering stage.
 Problem Student Response Diagnostic Checklist Verdict
-When $P=\sqrt{3}$ $-7\sqrt{5}+2\sqrt{3}$, $Q=2\sqrt{3}-\sqrt{5}$ $+2\sqrt{5}$, find the value of $P+Q$. [Uncaptioned image]
+When $P=\sqrt{3}$ $7\sqrt{5}+2\sqrt{3}$, $Q=2\sqrt{3}-\sqrt{5}$ $2\sqrt{5}$, find the value of $P+Q$. [Uncaptioned image]
 Recognize: Does the student
-realize that if the numbers in
-the radicals are the same, he
-or she can perform the four
-arithmetic operations on the
-coefficients with the same
-letters?
+realize that the coefficients can
+be combined when the radicals
+are the same?
 Vague
@@
 simply organize each equation
-(e.g., put $\sqrt{5}$ terms together,
-etc.) and then find the final
-sum?
+and then find the final sum?
 Evident
 No
-There is a trapezoid with a lower side length of $8$ $cm$ and a height of $4$ $cm$.
+There is a trapezoid with a lower side length of 88 cmcm and a height of 44 cmcm.
 If the area of this trapezoid is not less than $28$ $cm^{2}$, find how much more $cm$ the length of the upper side of the trapezoid must be.
@@
 remember the formula for
-the area of a shape correctly
-(e.g., the formula for the area
-of a trapezoid)?
+the area of a shape correctly?
 Evident
@@
 Table 1: Two samples from MathCog. Each data point is composed of a math problem, student response, relevant diagnostic checklist, and verdict for each check item.
+To evaluate LLMs’ performance on cognitive skill diagnosis, we created a benchmark dataset comprising secondary-school math problems, handwritten student responses, diagnostic checklists, and teacher-generated verdicts (Table 1).
+The math problem and student response data are from AI-Hub [^1] The data are available with approval from the provider via AIHub.
+, which provides OCR transcriptions of handwritten work.
+Each topic includes multiple isomorphic problems that share the same problem-solving procedures but differ in numerical values (e.g., “Solve $x^{2}+2x-3=0$” and “Solve $x^{2}+x-2=0$”) [25, 21].
+From this data, we focused on topics from grades 7–9, where problems are sufficiently complex to elicit students’ cognitive processes.
+We further excluded topics with fewer than 50 student responses to ensure sufficient data for reliable evaluation.
+We decided not to augment our dataset with synthesized data, as it might bias our observations on this novel task.
+Through this filtering process, the resulting dataset contains topics, 137 problems, and 796 student responses.
 3.1 Diagnostic Checklist
-For each problem type, we made a diagnostic checklist commonly applicable to isomorphic problems.
-These checklists are a list of binary question items mapped to one of the 15 cognitive skills defined in the TIMSS 2019 assessment framework (Mullis, 2017).
-The authors initially drafted the checklist items and refined them by requesting a review from five math curriculum and evaluation experts with PhD degrees in education and practical experience in making math assessment guidelines (e.g., TIMSS, Korean public school curriculum).
+For each topic, we developed a diagnostic checklist commonly applicable to isomorphic problems.
+Each checklist consists of binary question items mapped to one of the 15 cognitive skills defined in the TIMSS 2019 assessment framework [22].
+The checklist items were adapted from TIMSS skill descriptions and refined through expert review by five mathematics curriculum and evaluation experts with PhD degrees in education and practical experience in making math assessment guidelines.
 Experts gave feedback on the clarity, granularity, and validity of the checklist items.
-For example, the experts suggested changing “square root term” into “irrational term” for generality and combining items like “Were coefficients calculated correctly?” and “Were addition and subtraction performed correctly?” into “Were the addition and subtraction calculations of coefficients performed accurately?” for reasonable granularity: specific enough to describe a distinct skill, yet general enough to accommodate diverse forms of valid evidence.
 The experts also commented on the skills each problem can or cannot assess.
-Experts pointed out that our math problems primarily focus on calculating numbers and applying knowledge and, hence, are limited in assessing “reasoning” (e.g., justifying, analyzing, generalizing) (Mullis, 2017) by their design.
+Experts pointed out that our math problems primarily focus on calculating numbers and applying knowledge, and hence are limited in assessing “reasoning” (e.g., justifying, analyzing, generalizing) [22] by design.
 We scoped our check items to “knowing” and “applying” cognitive domains only.
@@
 We recruited 15 middle school math teachers to evaluate 796 student responses based on predefined diagnostic checklists.
-These teachers had an average of 6.1±4.3 years of experience (range: 2.5–20 years).
-Each check item was assessed along two dimensions: “yes/no” and “evident/vague” (see Table 1).
-A “yes” response indicated that the student fully demonstrated the cognitive action specified, while a “no” indicated otherwise.
+Teachers had an average of 6.1±4.3 years of experience (range: 2.5–20 years).
+Each check item was assessed along two dimensions: correctness (Yes/No) and evidential strength (Evident/Vague).
+A “Yes” response indicated that the student fully demonstrated the cognitive action specified, while a “No” indicated otherwise, including partially demonstrated responses.
 “Evident” meant there was clear evidence to support the judgment, whereas “Vague” signified insufficient evidence.
-For example, for the check item “Did the student calculate correctly?”, teachers marked “evident yes” if the response showed a complete, error-free calculation process, “evident no” if there were clear mistakes, and “vague no” if calculations were missing.
-To account for the subjectivity in the diagnosis, we grouped three teachers to evaluate approximately 160 student responses together.
-The diagnosis occurred in two phases: first, each group collaboratively assessed 36 responses, resolving conflicts through majority voting; second, each teacher independently assessed 18 responses.
-The second phase included 18 overlapping responses, allowing us to measure inter-rater agreement (see Table  4 in Appendix).
-Based on the threshold established in prior literature (Graham et al., 2012), we set a minimum agreement of 70% and excluded data from problem types 6, 13, and 14 accordingly.
-The distribution of the verdicts is presented in Table 5.
+Figure 1: Overview of the diagnostic pipeline of the MathCog.
+To account for the subjectivity in cognitive diagnosis, each student’s response was evaluated by three teachers with overlapping assignments, allowing us to measure inter-rater agreement (see Table  7 in Appendix [^2] Appendix can be found: OSF appendix materials).
+Following the threshold established in prior literature [7], we excluded topics with agreement below 70%, resulting in a final dataset of 12 topics, 110 problems, and 639 student responses.
+As a result, the benchmark focuses on diagnostic cases with stable expert agreement, providing a reliable basis for evaluating model performance.
+These topics cover three-fourths of the content domains defined in TIMSS 2019 (see Table 6 in Appendix), and the dataset remains reasonably sized given the substantial human effort required for expert annotation.
 4 Experimental Setting
-Using MathCog as a benchmark dataset, we explored two research questions.
-For RQ1, we evaluated the state-of-the-art LLMs, testing GPT (Achiam et al., 2023), Gemini (Team et al., 2023), Llama (Touvron et al., 2023), Claude (Anthropic, 2024), and DeepSeek (Guo et al., 2025) series.
-For RQ2, we configured three experiments in which we compared the performance of LLMs in different inputs, reasoning capabilities, and model sizes.
-Task Definition.
-Our task for LLMs is essentially a single-label classification task.
-Formally, we represent diagnosis result for a student’s response as a collection of tuples $(c,r)$, where $c\in C$ (i.e., check items) and $r\in\{Evident\ Yes,Vague\ Yes,Evident\ No,Vague\ No\}$.
-The combined judgment $(c,r)$ provides both an evaluative decision and a confidence signal for a skill.
-Given a context $(P,S,R,C)$, where P is the problem, S is the answer, and R is the student response, the task is to produce assessment tuples $(c,r)$ for all $c\in C$.
+Using MathCog as a benchmark dataset, we evaluated a diverse set of LLMs with varying input modalities, reasoning capabilities, model sizes, and prompting strategies to address RQ1, and analyzed how the evidential strength of student responses affects diagnosis performance to address RQ2.
 Prompting.
-We instructed LLMs to evaluate each check item based on a given math problem, its solution, a student’s response, and a diagnostic checklist.
-The LLMs processed inputs in the form of OCR transcriptions, where mathematical formulas and visual cues (e.g., strikethroughs) were represented using LaTeX.
-In the case of testing the effect of multimodality, we supplied images of student responses.
-Since the original inputs were in Korean, we machine-translated them into English to prevent possible performance degradation due to language (Achiam et al., 2023).
-To enhance response accuracy, we employed Chain-of-Thought prompting (Wei et al., 2022; Liu et al., 2024), guiding the LLM to systematically address each check item by first restating its content, identifying relevant evidence, providing an explanation, and delivering a final verdict.
+We instructed LLMs to evaluate each diagnostic check item given a math problem, its solution, a student’s response, and a diagnostic checklist.
+Student’s response image inputs were provided as OCR transcriptions, with mathematical formulas and visual cues (e.g., strikethroughs) represented in LaTeX.
+For multimodal settings, we additionally supplied images of student responses.
+Since the original inputs were in Korean, we machine-translated them into English to prevent possible performance degradation due to language [1]; we used Google Translate API for batch translation and manually verified them.
+We employed Chain-of-Thought prompting [16] to guide LLMs to systematically address each check item by first restating its content, identifying relevant evidence, providing an explanation, and delivering a final verdict.
 The verdict followed one of the four categories used by teachers in MathCog.
-We opted not to include few-shot examples, as MathCog does not provide explicit evidence and explanations for reference.
+To examine the effect of in-context examples on model performance, we experimented with few-shot prompting using two exemplars that covered all four types of verdicts.
 To minimize randomness, all LLMs were run with a temperature setting of zero.
-A detailed description of the system and user prompts can be found in the Appendix.
+Full system and user prompts are provided in the Appendix.
 Models.
-We conducted our experiment using 16 distinct closed and open LLMs from various vendors.
-To assess the impact of multimodal input, we tested GPT-4o (2024-08-06), Claude-3.5-Sonnet (20240620), and Gemini-1.5-Flash under two conditions: one using text-only prompts and another supplemented with an image.
-To examine the effect of reasoning capabilities, we compared LLMs with reasoning capabilities—GPT-o1-Preview (2024-09-12), Gemini-2.0-Flash-Thinking (exp-01-21), and DeepSeek-R1—with conventional LLMs of the same model series—GPT-4o, GPT-4o-Mini (2024-07-18), Gemini-1.5-Pro (002), Gemini-1.5-Flash, and DeepSeek-V3 (0324).
-Finally, we investigated model size differences: GPT-4o, Gemini-1.5-Pro, Llama-3.1-450B (2024-07-23), and DeepSeek-V3 were categorized as large models; Llama-3.3-70B (2024-07-23) and Qwen-2.5-72B (2024-09) as medium models; and GPT-4o-Mini, Gemini-1.5-Flash-8B (001), Llama-3.1-8B (2024-07-23), Qwen-2.5-7B (2024-09), and Mistral-7B (2024-08-22) as small models.
+We evaluated 18 LLMs from multiple model families, selected to analyze the effects of multimodality, reasoning capability, and model size.
+For multimodality, selected models were tested under both text-only and image-augmented conditions; for reasoning, we compared reasoning-oriented models with their conventional counterparts within the same model families.
 Metrics.
-We evaluated the outputs of the large language models (LLMs) against ground-truth diagnostic results provided by human experts.
-Performance was assessed using both macro F1 score and accuracy computed over verdicts (i.e., $(c,r)$) to quantify the effectiveness of skill diagnosis.
-For a more nuanced analysis, we also examined the models’ tendencies toward overconfidence and underconfidence.
-Specifically, overconfidence is defined as the conditional proportion of “evident” verdicts produced by the LLMs that were incorrect.
-Conversely, underconfidence is the conditional proportion of verdicts labeled as “vague” by the LLMs that were, in fact, both correct and should have been classified as “evident.”
+We evaluated LLM outputs against teacher-provided ground-truth diagnostic labels.
+Overall diagnosis performance was assessed using macro F1 score and accuracy over the four verdict categories.
+We report macro F1 to mitigate the effect of label imbalance across verdicts (see Table 8 in Appendix).
+To address RQ2, we analyzed how the evidential strength of student responses affects diagnosis performance.
+To formalize this analysis, we treat each diagnosis instance as a $(\textit{student response},\textit{cognitive skill})$ pair, indexed by $i$.
+For each instance $i$, the ground-truth labels consist of a skill judgment $y_{i}\in\{\text{Yes},\text{No}\}$ and evidential strength $e_{i}\in\{\text{Evident},\text{Vague}\}$, and the corresponding model predictions $\hat{y}_{i}$ and $\hat{e}_{i}$.
+To characterize how models handle evidential strength, we define two complementary metrics that capture distinct failure modes in evidence attribution.
+First, we define evidence over-attribution (OverAttr) as the frequency with which a model assigns “Evident” to cases where the ground-truth evidential strength is “Vague”:
+$\text{OverAttr}=\frac{\left|\left\{i\mid e_{i}=\text{Vague}\wedge\hat{e}_{i}=\text{Evident}\right\}\right|}{\left|\left\{i\mid e_{i}=\text{Vague}\right\}\right|}$ (1)
+This metric reflects the model’s tendency to overstate evidential support under vague evidence cases, independent of diagnosis correctness.
+Second, we define evidence false-attribution (FalseAttr) as the frequency with which a model assigns Evident to incorrect diagnoses:
+$\text{FalseAttr}=\frac{\left|\left\{i\mid\hat{y}_{i}\neq y_{i}\wedge\hat{e}_{i}=\text{Evident}\right\}\right|}{\left|\left\{i\mid\hat{y}_{i}\neq y_{i}\right\}\right|}$ (2)
+This metric captures a particularly misleading failure mode, where erroneous diagnoses are accompanied by strong evidential claims.
 5 Results
-In this section, we organize and interpret the findings to answer our research questions. The complete results from our experiment are presented in Table 6 in the Appendix.
-5.1 RQ1. How accurately do state-of-the-art LLMs diagnose students’ cognitive skills?
-Overall, the results underscore the difficulty of the task (Table 2); no model surpassed an F1 score of 0.5, indicating that even state-of-the-art models struggle to diagnose students’ cognitive skills accurately.
-Our manual error analysis on a few samples revealed that LLMs largely failed when 1) student response flows in a non-directional, scattered layout, 2) a student uses underscores, strikethroughs, and calculation notes, and 3) a response lacks direct evidence for a skill, despite contextual information being sufficient for inference.
-[…]
-Models Precision Recall F1 Accuracy Overconfidence Underconfidence
-DeepSeek-V3 .406 .461 .417 .714 .818 .107
-GPT-4o .398 .476 .412 .672 .788 .110
-Llama-3.1-405B .415 .438 .384 .624 .778 .186
-Claude-3.5-Sonnet .329 .387 .332 .654 .812 .173
-Gemini-1.5-Pro .430 .496 .441 .707 .840 .132
-Table 2: The performance of the five state-of-the-art LLMs on the cognitive skill diagnosis task.
+This section reports results addressing our research questions.
+Category Model Precision Recall F1 Accuracy OverAttr FalseAttr
+Reasoning Model DeepSeek-R1 .447 .467 .442 .773 .767 .862
+Gemini-2.0-Flash-Thinking .454 .486 .418 .688 .361 .432
+o1-Preview .423 .513 .443 .711 .489 .650
+Multimodal Model Claude-3.5-Sonnet-img .408 .457 .413 .691 .626 .559
+Gemini-1.5-Flash-img .431 .484 .429 .702 .471 .471
+GPT-4o-img .427 .494 .448 .743 .656 .681
+Large Model Claude-3.5-Sonnet .411 .487 .416 .656 .551 .574
+DeepSeek-V3 .406 .461 .417 .714 .612 .595
+Gemini-1.5-Pro .429 .496 .440 .706 .520 .635
+GPT-4o .398 .476 .412 .672 .648 .658
+Llama-3.1-405B .415 .439 .385 .624 .467 .307
+Medium Model Llama-3.3-70B .415 .467 .397 .635 .436 .409
+Qwen-2.5-72B .415 .459 .416 .711 .568 .500
+Small Model Gemini-1.5-Flash .432 .490 .432 .679 .502 .465
+Gemini-1.5-Flash-8b .359 .427 .348 .558 .537 .525
+GPT-4o-mini .345 .398 .352 .622 .767 .784
+Llama-3.1-8B-128K .352 .334 .323 .653 .709 .624
+Qwen-2.5-7B .347 .337 .339 .705 .758 .798
+Table 2: Performance of all 18 LLMs tested on skill diagnosis tasks. For each metric, blue and red fonts indicate the best and worst values, respectively.
+Figure 2: F1 score across by model type and family.
+5.1 RQ1. How do different LLMs perform in diagnosing students’ cognitive skills?
+Overall, all evaluated LLMs showed limited performance in cognitive skill diagnosis, with macro F1 scores below 0.5 (Figure 2, Table 2).
+While overall accuracy was relatively moderate ($M=.680$, $SD=.050$), it obscures important diagnostic failures.
+Models frequently over-attributed evidential strength (OverAttr; $M=.580$, $SD=.120$) and falsely attributed evidence (FalseAttr; $M=.585$, $SD=.145$), often asserting strong evidence even when diagnoses were incorrect.
+Notably, even the best-performing models in terms of F1 score and accuracy (e.g., DeepSeek-R1, GPT-4o-img) showed high OverAttr and FalseAttr values, suggesting that performance gains often co-occur with unreliable evidential judgement.
+Figure 3: Performance and evidential attribution errors across cognitive skills, measured by F1 score, accuracy, evidence over-attribution, and false-attribution.
 Analysis of skill-specific performance (Figure 3) reveals further insights into the limitations of current LLMs.
-No skill category achieved a maximum F1 score above 0.5.
-Interestingly, the skill Compute, which most closely resembles conventional grading tasks (i.e., evaluating the correctness of mathematical procedures and calculations), still yielded only a moderate F1 score across models.
-This moderate F1 score confirms LLMs’ struggle in verifying correct computation in constructed student responses.
-Contrary to expectations, there was no substantial performance gap between the “Knowing” and “Applying” cognitive skills.
-Although we anticipated that “Knowing” skills (e.g., Recall, Recognize) would be easier than “Applying” skills for LLMs due to their surface-level nature, models performed comparably.
-An exception is found in the skill Classify/Order, which shows a relatively high average accuracy but a lower F1 score.
-Closer inspection reveals that “Evident Yes” is dominant in the ground truth dataset (Table 5), leading to the discrepancy between F1 score and accuracy.
-Figure 2: The performance of five state-of-the-art LLMs across cognitive skill types.
-The left plot shows the minimum (bottom $\times$), maximum (top $\times$), and average (middle •) F1 scores for each skill category, while the right plot presents the corresponding accuracy metrics.
-5.2 RQ2. How do multimodal input, reasoning, and model size influence LLMs’ cognitive skill diagnosis?
-We observed subtle impacts of multimodal input and reasoning on the performance, while the impact of model size was significant.
-Multimodal Input.
-Models supporting image input (e.g., GPT-4o-img, Claude-3.5-Sonnet-img, Gemini-1.5-Flash-img) consistently outperformed their text-only counterparts in accuracy (up to 10.6%) and in reducing underconfidence (up to 4.9%) (Figure 3 left), although the statistical difference between the pairs was not significant.
-Image input was particularly beneficial in cases where OCR transcription introduced errors.
-For example, a student underlined an arithmetic expression and wrote the corresponding calculation beneath it (see Figure 5 a).
-The OCR system misinterpreted this layout as a fraction, leading Claude-3.5-Sonnet (text-only) to incorrectly conclude that the student had set up an invalid equation.
-In contrast, the same model with image input correctly inferred the student’s intention and produced an accurate diagnosis.
-Reasoning.
-The performance of reasoning-oriented models showed mixed results (Figure 3 right).
-DeepSeek-R1 achieved the highest F1 and accuracy scores among all models.
-However, other reasoning models, such as o1-Preview and Gemini-2.0-Flash-Thinking, did not show clear advantages over standard models in either accuracy or F1 score.
-These results suggest that reasoning capability alone does not guarantee better performance.
-For instance (Figure 5 b), DeepSeek-R1 classified a response as “Vague No” because the student omitted the factorization steps.
-However, the original expression was $\sqrt{48a}$, and the model failed to detect a clear error in the variable term, indicating that the student’s reasoning was evidently flawed.
-Throughout its reasoning process, the model focused on the constant term only and overlooked other possible places of error.
-While current reasoning models excel at pushing forward linear, step-by-step logic, effective cognitive diagnosis may also require divergent, multi-perspective reasoning for thorough analysis.
-[…]
-Figure 3: Impact of multimodal input (left) and reasoning capability (right) on LLMs’ performance in cognitive skill diagnosis.
-On the left, each adjacent pair of bars shows performance with and without the image input from the same model family.
-On the right, models with reasoning capabilities are grouped at the front.
-Figure 4: Impact of model size on LLM performance in cognitive skill diagnosis. Models are grouped into large, medium, and small categories.
+No skill category achieved an F1 score above 0.5, indicating that fine-grained diagnosis remains challenging even at the individual skill level.
+Despite this overall limitation, we observed a performance gap between the Knowing and Applying cognitive skills.
+Contrary to our expectation that Knowing skills (e.g., Recall, Recognize) would be easier due to their surface-level nature, models performed better on Applying skills in both F1 score ($t=4.34$, $p<.001$) and accuracy ($t=4.85$, $p<.001$).
+However, this performance gain came with increased evidential errors.
+Applying skills exhibited higher evidence over-attribution ($t=5.44$, $p<.001$) and false-attribution ($t=7.66$, $p<.001$) than Knowing skills, indicating a tendency to overstate evidential support when diagnosing higher-level cognitive skills.
+Few-shots Precision Recall F1 Accuracy OverAttr FalseAttr
+claude-3-5-sonnet .398 .485 .410 .646 .515 .647
+deepseek-V3 .404 .493 .404 .624 .493 .497
+gemini-1.5-pro .429 .517 .430 .665 .427 .552
+gpt-4o .390 .485 .396 .637 .577 .587
+llama-3.1-405B .442 .494 .429 .677 .476 .457
+Table 3: Performance of five state-of-the-art models in few-shot experiments. Bold values denote improvements over the corresponding zero-shot results.
+We further examined the impact of multimodality, reasoning, model size, and few-shot prompting as factors that may influence diagnostic performance (Table  2).
+Multimodal input provided modest but inconsistent benefits.
+Models supporting image input consistently outperformed their text-only counterparts in accuracy, indicating that access to handwritten layouts and visual cues helps mitigate errors introduced by OCR transcription.
+However, improvements in F1 score were inconsistent, suggesting that visual input can sometimes introduce noise and does not uniformly improve fine-grained skill diagnosis.
+Reasoning-oriented models produced similarly mixed results.
+While some reasoning-oriented models (e.g., DeepSeek-R1, GPT-o1) achieved higher F1 scores and accuracy than their non-reasoning counterparts, Gemini performed worse.
+These results suggest that reasoning mechanisms do not uniformly translate into improved cognitive diagnosis.
+Model size showed a moderate positive relationship with performance.
+Larger models generally achieved higher F1 scores than smaller ones, indicating that increased parameter capacity supports a more robust interpretation of student responses.
+However, the correlation was moderate (Spearman’s $\rho=.570,p=.053$), suggesting that scale alone is insufficient to account for the substantial variability in performance across models.
+Finally, few-shot prompting did not improve diagnostic performance.
+While in-context examples might seem helpful for such nuanced tasks, F1 and accuracy often decreased rather than improved (Table 3), suggesting the challenge lies in the complexity of cognitive skill diagnosis rather than prompt design.
+However, few-shot prompting did reduce tendencies toward evidence over-attribution and false-attribution, indicating that examples may encourage more cautious, evidence-grounded responses even when overall diagnostic accuracy remains limited.
+Taken together, these findings suggest that prompt engineering and model capabilities have limited impact on improving the performance of cognitive skill diagnosis.
+5.2 RQ2. How does the evidential strength of student responses affect LLM’s cognitive skill diagnosis performance?
+Figure 4: Effect of evidential strength on LLM diagnostic performance for every model.
+Our results reveal a strong dependency of LLM diagnostic performance on the evidential strength of student responses.
+Across all models, both F1 score ($t=64.60$, $p<.001$) and accuracy ($t=17.80$, $p<.001$) were substantially higher when student responses contained evident evidence for a cognitive skill than when evidence was vague (Fig.
+4).
+The magnitude of this gap was substantial (mean $\Delta F1=.51$; mean $\Delta Acc=.41$), indicating that current LLMs rely heavily on explicit evidence and exhibit limited robustness when reasoning under weak or implicit evidential conditions.
+Beyond degraded performance, models also exhibited over-attribution bias under vague evidence.
+Specifically, they frequently labeled responses as Evident even when the ground-truth evidential strength was Vague, resulting in consistently high OverAttr values.
+More concerningly, models also showed a strong tendency toward FalseAttr, in which they asserted evident evidence despite arriving at incorrect cognitive skill diagnoses.
+In such cases, models produced unsupported explanations that appeared plausible yet were not grounded in students’ actual responses.
+Notably, these two tendencies were strongly correlated across models (Spearman’s $\rho=.837$, $p<.001$), indicating that models prone to over-attributing evidential strength are also more likely to produce incorrect diagnoses with asserted evidence.
+This behavior is particularly problematic for educational use, as it can mislead teachers and students by presenting incorrect diagnoses with seemingly well-justified rationales [15].
+Error code Description Count (%)
+E1. Evidence Misidentification Incorrectly identifying or relying on insufficient evidence from the student response. 659 (23.40%)
+E2. Rubric Misinterpretation Misunderstanding the diagnostic rubric or evaluation criteria. 279 (9.91%)
+E3. Over-Inference Inferring unstated reasoning beyond the information provided in the student response. 942 (33.45%)
+E4. Consistency Issue Inconsistencies between evidence, explanation, and final verdict. 379 (13.46%)
+E5. Hallucination Introducing content or reasoning not present in the student response. 557 (19.78%)
+Table 4: Error types and their distribution identified from qualitative analysis of LLM diagnostic failures under vague evidence.
+To better understand how such errors arise, we conducted a qualitative error analysis of model outputs under incorrectly diagnosed vague evidence cases.
+We examined all components generated in the models’ CoT responses, including the identified evidence, explanations, and final verdicts, focusing on how the models misinterpreted students’ answers, generated unsupported reasoning, and ultimately reached incorrect diagnoses.
+Three authors collaboratively analyzed 372 cases (13.2%) sampled from a total of 2,816 error instances, covering 18 model responses, and iteratively derived five recurring error types with substantial inter-rater reliability (Fleiss’ $\kappa=0.74$).
+The remaining cases were then independently labeled using the agreed error taxonomy.
+Table  4 summarizes the distribution of error types across model categories.
+Errors related to over-inference were most prevalent (33.5%), followed by evidence misidentification (23.4%) and hallucination (19.8%).
+We further observed systematic differences across model types (Figure  5): smaller models exhibited a higher proportion of hallucination errors (48.78%), whereas reasoning-oriented models committed over-inferred more frequently (46.61%) beyond the information explicitly provided in student responses.
+Figure 5: Error type distribution by model category.
 6 Discussion
-Our findings imply that current LLMs are not yet suitable for high-stakes cognitive skill diagnosis.
-The models exhibit low recall and high overconfidence (Table 2), raising concerns that LLM-based systems may frequently overlook students’ misconceptions.
-This shortcoming is particularly problematic given that the primary goal of automated assessment is to identify and address such misconceptions, and it can compromise teachers’ and students’ trust in automated systems (Shin, 2021).
-To responsibly integrate LLMs into real-world assessment settings, practitioners should begin by evaluating their performance on small, ground-truth samples.
-This process not only exposes limitations but also allows for fine-tuning model judgments and confidence levels for downstream math topics and student populations.
-In parallel, model developers may work to mitigate overconfidence during training by encouraging more conservative, evidence-based decision-making.
-Our manual analysis of failed cases reveals that inferring students’ thought processes from handwritten responses presents distinct challenges.
-While mathematical problem-solving is inherently procedural, students often express their reasoning in non-linear, spatially scattered layouts that lack clear directionality.
-This ill-structured presentation differs significantly from the format of most existing reasoning and problem-solving benchmarks.
-Moreover, ambiguous notations (e.g., fractions misinterpreted as underscores) and unconventional layouts frequently result in OCR transcription errors, which hinder accurate diagnosis unless image inputs are provided.
-To enable more reliable inference, future work may require richer representations of math responses that capture high-level semantic intent beyond surface-level positional and textual cues.
-7 Limitations and Future Work
-While MathCogrepresents a novel benchmark for evaluating LLMs on cognitive skill diagnosis, it has several limitations.
-First, due to data-sharing restrictions from the source provider (AIHub), the original math problems and student responses cannot be publicly released outside Korea.
-To support future research despite this constraint, we release the complete set of diagnostic checklists and expert-generated human annotations, which can serve as references for researchers to develop similar datasets in their educational contexts.
-In parallel, we are actively collecting new student response data under informed consent, with the goal of constructing a fully public benchmark in the future.
-This ongoing data collection effort will also address limitations in the diversity of problem types represented in MathCog.
-Currently, the dataset focuses on nine problem types, primarily centered around arithmetic equation solving.
-Other areas of mathematics—such as geometry—often require interpreting visual representations and may further benefit from the integration of multimodal input.
-Additionally, reasoning-oriented problems, which explicitly ask students to generalize mathematical knowledge or justify their arguments, are underrepresented.
-We plan to expand the dataset to include a broader range of content areas and grade levels, enabling a more holistic evaluation of LLM performance across the mathematics curriculum.
-Our experimental results are also bound to zero-shot settings, without exploring the effects of few-shot or test-time compute (Snell et al., 2024) prompting.
-Future work can investigate whether more sophisticated prompting enhances diagnostic accuracy.
-For example, prompting an LLM to explicitly infer a student’s thinking process before answering each check item may lead to more accurate judgments.
-We see our work as a first step toward automated cognitive skill assessment, and we invite the community to build upon these directions.
-Acknowledgments
-This work was supported by Algorithm LABS.
-This research (paper) used datasets from “The Open AI Dataset Project (AI-Hub, S.
-Korea).” All data information can be accessed through “AI-Hub (www.aihub.or.kr).”
-Appendix A Appendix
+Our findings imply that current LLMs are not yet reliable for cognitive skill diagnosis without intermediate verification, as errors cascade throughout the diagnostic pipeline.
+Performance was strongly constrained by the evidential strength in student responses rather than by model capabilities.
+This challenge is closely tied to the nature of students’ handwritten work, which often omits intermediate reasoning and expresses ideas in implicit or non-linear ways, making diagnostic evidence difficult to assess [11, 24].
+Under such conditions, LLMs tend to over-attribute evidential strength and, in some cases, arrive at incorrect diagnoses with asserted evidence, raising concerns for real-world educational use.
+Our findings point to an important opportunity: richer diagnostic evidence could improve cognitive diagnosis.
+One promising direction is to design problem-solving tasks and interfaces that better elicit students’ reasoning processes and justifications, help them organize their steps tidily, and prompt reflection on strategy choices.
+For example, interactive interfaces could track diagnostic evidence in real-time and request clarification when evidence appears weak, shifting from post-hoc inference to active elicitation of cognitive processes.
+Beyond task design, our qualitative error analysis highlights the importance of teacher-in-the-loop approaches for mitigating diagnostic failures.
+Diagnostic errors followed recurring patterns across key stages of the diagnostic process, including identifying evidence, interpreting diagnostic criteria, and synthesizing intermediate reasoning into final judgments.
+At these stages, teachers can intervene in targeted ways.
+Errors such as evidence misidentification can be mitigated by having teachers validate whether the identified evidence is actually present in the student’s work.
+Rubric misinterpretation can be addressed by allowing teachers to check and correct how model interpretations align with diagnostic criteria.
+For over-inference, teachers can scrutinize whether inferred cognitive skills are sufficiently supported by observable evidence, preventing unwarranted extrapolation.
+Consistency issues and hallucinations require human oversight to ensure coherence between intermediate reasoning and final diagnostic judgments.
+Rather than treating model outputs as final assessments, such selective human intervention can reduce diagnostic errors, consistent with prior work on evidence-supported reasoning [28] and human-AI complementary assessment workflows [26].
+Overall, responsibly integrating LLMs into educational assessment requires moving beyond fully automated diagnosis toward transparency and selective human verification.
+7 Future Work
+While we present a novel, carefully crafted benchmark for evaluating LLM-based cognitive skill diagnosis, several directions remain for future work.
+First, the benchmark can be extended in a more targeted manner.
+Synthetic data generation [3] could selectively augment diagnostically challenging cases, such as student responses with implicit or vague evidence.
+Our analysis of 3,036 verdicts highlights where such challenges arise, motivating controlled data expansion rather than indiscriminate scaling.
+Second, future work can broaden the range of cognitive demands represented in the dataset.
+Reasoning-oriented problems that require students to generalize or justify their thinking are currently underrepresented in MathCog, and expanding coverage across problem types, topics, and grade levels would enable more comprehensive evaluation.
+Finally, beyond dataset expansion, future work can explore richer evaluation settings for cognitive diagnosis, such as human–AI collaborative diagnosis scenarios or analyzing how LLMs revise diagnostic judgments when provided with additional evidence.
+It could shed light on the reliability and practical use of LLM-based diagnostic systems in educational contexts.
+Appendix 0.A Appendix
+0.A.1 Prompts
+The blue text represents the programmatically filled arguments, and the orange text represents LLM-generated output.
+0.A.2 System Prompt
+# **Task Description**
+You are a middle school math teacher tasked with evaluating students mathematical thinking skills based on their responses to math problems.
+Your goal is to analyze a given students response and determine whether they exhibit specific cognitive skills in solving the problem.
+Your evaluation must be **strict** and **evidence-based**, meaning that every verdict must be backed by direct evidence from the response.
+If no clear evidence exists, do not assume correctness.
+## **Evaluation Categories**
+For each thinking skill in the checklist, you must classify the students performance into one of the following categories:
+- **Evident Yes**: The students response provides clear and explicit evidence that the check item is met. A direct quote from the response can confirm this.
+- **Vague Yes**: The response suggests that the check item might be satisfied, but no specific part of the response directly proves it.
+- **Evident No**: The response explicitly contradicts or fails to meet the check item, with clear evidence demonstrating the error or omission.
+- **Vague No**: The response does not appear to satisfy the check item, but there is no direct evidence confirming whether the student considered it or not.
+## **Input Format**
+You will receive the following data:
+- **Problem**: A math problem given to a student.
+- **Answer**: The correct step-by-step solution.
+- **Response**: The students response to the problem.
+- **Check Items**: A set of specific skills to evaluate.
+## **Output Format**
+Return a valid JSON object structured as follows:
+‘‘‘json
+"skills": [
+"checkItem": "<Check Items [Label] and the Following Question>",
+"evidence": "<Directly Quoted Part of Response>",
+"explanation": "<Explanation About Why the Evidence Supports the Verdict>",
+"verdict": "Evident Yes" | "Vague Yes" | "Evident No" | "Vague No"
+0.A.3 User Prompt
+# **Task**
+student responses to math problems, extract direct evidence, and strictly classify thinking skills according to the given categories.
+Return a valid JSON object structured as follows:
+‘‘‘json
+"skills": [
+"checkItem": "<Check Items [Label] and the Following Question>",
+"evidence": "<Directly Quoted Part of Response>",
+"explanation": "<Explanation About Why the Evidence Supports the Verdict>",
+"verdict": "Evident Yes" | "Vague Yes" | "Evident No" | "Vague No"
+**Problem:**
+There is a two-digit natural number whose tens digit is 11. If the number that changes the tens and ones digits of this natural number is 99 less than 55 times the first number, find the first number.
+**Answer:**
+If the number in the ones place is $x$, this natural number is $10+x$.
+The number where the tens digit and the ones digit are swapped is $10x+1$ because the tens digit is $x$ and the ones digit is 11.
+The changed number is 99 smaller than 55 times the first number, so 10x+1=5(10+x)910x+1=5(10+x)-9 $10x+1=5x+41$ $5x=40$ $x=8$ Therefore, the first number is $18$.
+**Response:**
+11 $x$ $x$ 11 10x+1=5(10+x)910x+1=5(10+x)-9 10x+1=50+5x910x+1=50+5x-9 $41$ $5x=40$ $x=8$
+**Check Items:**
+- [recognize_0] Is the student aware that the relationship between two numbers switched can be expressed as addition/subtraction/multiplication?
+- [recognize_1] Has the student figured out all the elements necessary to solve the problem (original numbers, swapped numbers, difference or multiple relationships, etc.)?
+- [compute_0] Has the student performed algebraic calculations correctly, such as expanding expressions and organizing similar terms?
+- [determine_0] Has the student chosen the appropriate approach to solving the problem (e.g., establishing a linear equation)?
+- [represent_0] Are two-digit numbers expressed correctly in the form 10x + y (or equivalent)?
+- [implement_0] Does the student proceed consistently with the process of expanding the equation and moving terms to find x (or y) according to the planned solution strategy?
+**Output:**
+"skills": [
+"checkItem": "[recognize_0] Is the student aware that the relationship between two numbers switched can be expressed as addition/subtraction/multiplication?",
+"evidence": "10 x+1=5(10+x)-9",
+"explanation": "The student sets up the equation 10x+1=5(10+x)-9, which shows they understand the relationship between the original and swapped numbers involves subtraction and multiplication.",
+"verdict": "Evident Yes"
+"checkItem": "[recognize_1] Has the student figured out all the elements necessary to solve the problem (original numbers, swapped numbers, difference or multiple relationships, etc.)?",
+"evidence": "10 x+1=5(10+x)-9",
+"explanation": "The student uses the correct equation to represent the relationship between the original and swapped numbers, indicating they have identified the necessary elements.",
+"verdict": "Evident Yes"
+0.A.4 MathCog Dataset Details
+This section provides supplementary details of the MathCog dataset, including the diagnostic workflow, cognitive skill taxonomy, inter-rater reliability of teachers’ diagnoses, dataset composition, and representative failure cases of model diagnoses.
 Domains Cognitive Skills Description
@@
 Compute
-Carry out algorithmic procedures for +,,×,÷+,-$\times$,$\div$+ , - , × , ÷ or a combination of these
+Carry out algorithmic procedures for $+,-,\times,\div$ or a combination of these
 with whole numbers, fractions, decimals, and integers. Carry out
@@
 Table 5: Fifteen cognitive skills and their descriptions defined in the TIMSS 2019 framework.
-Problem Types 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
+Topics Content Domain Difficulty
+1. Problems involving the digits of numbers Number 1, 3
+2. Solving for unknowns under special conditions on the solution Algebra 1, 2, 3
+3. Finding unknowns when two equations have the same solution Algebra 1, 2
+4. Applying linear inequalities to geometric figures Geometry 2, 3
+5. Applying linear inequalities to pricing Algebra 1, 2
+7. Using square roots to calculate lengths in geometric figures Geometry 2, 3
+8. Performing simple addition and subtraction of square roots Number 1, 2
+9. Applying multiplication formulas Number 2, 3
+10. Determining coefficients or constants to complete the square Number 1, 2, 3
+11. Solving quadratic equations by factoring Algebra 1, 2
+12. Rewriting expressions in perfect square form Algebra 1, 2
+15. Finding the quadratic function given the vertex and another point on its graph Algebra 1, 2
+Table 6: The topics and difficulty levels of problems.
+This information comes from the original dataset’s metadata, which details the topic and difficulty level (1-3) of each problem.
+The isomorphic problems require the same mathematical concept to solve, but the difference in numbers makes one more tricky and complicated than the other.
+Topics 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
 % Agreement 95 70 96 95 89 68 74 86 80 80 100 88 40 57 78
-Table 4: The inter-rater percentage absolute agreement of each problem type. The percentage indicates the ratio of unanimous verdicts in each teacher group.
-A.1 Prompts
-The blue text represents the programmatically filled arguments, and the orange text represents LLM-generated output.
-A.2 System Prompt
-# **Task Description**
-You are a middle school math teacher tasked with evaluating students mathematical thinking skills based on their responses to math problems.
-Your goal is to analyze a given students response and determine whether they exhibit specific cognitive skills in solving the problem.
-Your evaluation must be **strict** and **evidence-based**, meaning that every verdict must be backed by direct evidence from the response.
-If no clear evidence exists, do not assume correctness.
-## **Evaluation Categories**
-For each thinking skill in the checklist, you must classify the students performance into one of the following categories:
-- **Evident Yes**: The students response provides clear and explicit evidence that the check item is met. A direct quote from the response can confirm this.
-- **Vague Yes**: The response suggests that the check item might be satisfied, but no specific part of the response directly proves it.
-- **Evident No**: The response explicitly contradicts or fails to meet the check item, with clear evidence demonstrating the error or omission.
-- **Vague No**: The response does not appear to satisfy the check item, but there is no direct evidence confirming whether the student considered it or not.
-## **Input Format**
-You will receive the following data:
-- **Problem**: A math problem given to a student.
-- **Answer**: The correct step-by-step solution.
-- **Response**: The students response to the problem.
-- **Check Items**: A set of specific skills to evaluate.
-## **Output Format**
-Return a valid JSON object structured as follows:
-‘‘‘json
-"skills": [
-"checkItem": "<Check Items [Label] and the Following Question>",
-"evidence": "<Directly Quoted Part of Response>",
-"explanation": "<Explanation About Why the Evidence Supports the Verdict>",
-"verdict": "Evident Yes" | "Vague Yes" | "Evident No" | "Vague No"
-A.3 User Prompt
-# **Task**
-student responses to math problems, extract direct evidence, and strictly classify thinking skills according to the given categories.
-Return a valid JSON object structured as follows:
-‘‘‘json
-"skills": [
-"checkItem": "<Check Items [Label] and the Following Question>",
-"evidence": "<Directly Quoted Part of Response>",
-"explanation": "<Explanation About Why the Evidence Supports the Verdict>",
-"verdict": "Evident Yes" | "Vague Yes" | "Evident No" | "Vague No"
-**Problem:**
-There is a two-digit natural number whose tens digit is $1$.
-If the number that changes the tens and ones digits of this natural number is $9$ less than $5$ times the first number, find the first number.
-**Answer:**
-If the number in the ones place is $x$, this natural number is $10+x$.
-The number where the tens digit and the ones digit are swapped is $110x+$ 1 because the tens digit is $x$ and the ones digit is $1$.
-The changed number is $9$ smaller than $5$ times the first number, so $910x+1=5(10+x)-$ 9 $4110x+1=5x+$ 41 $405x=$ 40 $x=8$ Therefore, the first number is $18$.
-**Response:**
-$1$ $x$ $x$ $1$ $910x+1=5(10+x)-$ 9 $910x+1=50+5x-$ 9 $41$ $405x=$ 40 $x=8$
-**Check Items:**
-- [recognize_0] Is the student aware that the relationship between two numbers switched can be expressed as addition/subtraction/multiplication?
-- [recognize_1] Has the student figured out all the elements necessary to solve the problem (original numbers, swapped numbers, difference or multiple relationships, etc.)?
-- [compute_0] Has the student performed algebraic calculations correctly, such as expanding expressions and organizing similar terms?
-- [determine_0] Has the student chosen the appropriate approach to solving the problem (e.g., establishing a linear equation)?
-- [represent_0] Are two-digit numbers expressed correctly in the form 10x + y (or equivalent)?
-- [implement_0] Does the student proceed consistently with the process of expanding the equation and moving terms to find x (or y) according to the planned solution strategy?
-**Output:**
-‘‘‘json
-"skills": [
-"checkItem": "[recognize_0] Is the student aware that the relationship between two numbers switched can be expressed as addition/subtraction/multiplication?",
-"evidence": "10 x+1=5(10+x)-9",
-"explanation": "The student sets up the equation 10x+1=5(10+x)-9, which shows they understand the relationship between the original and swapped numbers involves subtraction and multiplication.",
-"verdict": "Evident Yes"
-"checkItem": "[recognize_1] Has the student figured out all the elements necessary to solve the problem (original numbers, swapped numbers, difference or multiple relationships, etc.)?",
-"evidence": "10 x+1=5(10+x)-9",
-"explanation": "The student uses the correct equation to represent the relationship between the original and swapped numbers, indicating they have identified the necessary elements.",
-"verdict": "Evident Yes"
-Recall Recognize Classify/Order Compute Retrieve Determine Represent Implement
-Evident Yes 507 506 92 584 9 441 154 319
-Evident No 78 69 8 160 9 54 55 146
-Vague Yes 58 54 2 48 0 25 6 29
-Vague No 98 62 0 58 34 16 46 41
-Student responses 582 583 51 796 52 482 261 535
+Table 7: The inter-rater percentage absolute agreement of each topic. The percentage indicates the ratio of unanimous verdicts in each teacher group.
+Recall Recognize Classify/Order Compute Determine Represent Implement
+Evident Yes 413 434 92 522 441 122 307
+Evident No 62 56 8 125 54 33 140
+Vague Yes 20 16 2 26 25 0 18
+Vague No 35 28 0 20 16 3 18
+Student responses 530 534 102 693 536 158 483
 Table 8: Distribution of verdicts and number of diagnosed student responses in each cognitive skill.
 Note that the number of student responses can be larger than the sum of the four labels because some diagnostic checklists have two items for the same cognitive skill.
-Models Precision Recall F1 Accuracy Overconfidence Underconfidence
-DeepSeek-R1 .447 .467 .442 .773 .901 .028
-DeepSeek-V3 .406 .461 .417 .714 .710 .107
-GPT-4o-img .427 .494 .448 .743 .738 .070
-GPT-4o .398 .476 .412 .672 .703 .110
-GPT-4o-Mini .345 .397 .352 .622 .791 .102
-o1-Preview .338 .410 .354 .711 .757 .117
-Llama-3.3-70B .416 .468 .400 .640 .534 .189
-Llama-3.1-8B .283 .265 .257 .642 .695 .138
-Llama-3.1-405B .415 .438 .384 .624 .451 .186
-Claude-3.5-Sonnet-img .326 .365 .330 .690 .680 .140
-Claude-3.5-Sonnet .329 .387 .332 .654 .676 .173
-Gemini-1.5-Flash-img .431 .484 .429 .702 .593 .129
-Gemini-1.5-Flash-8b .287 .341 .279 .558 .579 .196
-Gemini-1.5-Flash .432 .490 .432 .679 .582 .155
-Gemini-1.5-Pro .430 .496 .441 .707 .721 .132
-Gemini-2.0-Flash-Thinking .455 .492 .421 .690 .589 .186
-Mistral-7B .384 .369 .347 .714 .805 .112
-Qwen-2.5-7B .278 .270 .271 .705 .842 .075
-Qwen-2.5-72B .415 .459 .416 .710 .631 .115
-Table 6: Performance comparison of all 16 LLMs tested on skill diagnosis tasks.
 Figure 6: Illustrative examples of diagnosis check items and student responses that LLMs failed to diagnose correctly. Evidence for human judgment is marked with a red box.